# Channel-Level Calibration Methods of Silicon Photomultiplier for JUNO-TAO Central Detector

### v1

@@ 5.1 Gain
The gain is quantified as the average charge output per single photoelectron (PE) avalanche in the SiPM after amplification by the front-end electronic board [Ref29]. Consequently, the gain depends on the operating voltages of both the SiPM and the front-end board. In the TAO CD, these two voltages are stable.
$\begin{split}f=\sum_{i=1}^{n}N_{i}\cdot Gaus(\mu_{i},\sigma_{i})\end{split}$ (5.1)
@@
As shown in Figure 8, the charge generated by a single PE follows a Gaussian distribution, and the individual photons detected by the avalanche photodiodes (APDs) in the SiPM are independent. Consequently, the total charge resulting from the simultaneous detection of multiple PEs also follows a Gaussian distribution, and the overall distribution is a multi-Gaussian distribution (Equation 5.1 where f is the multi-Gaussian function, $N_{i}$ is the number of hits corresponding to each PE, $\mu_{i}$ is the mean value, and $\sigma_{i}$ is the standard deviation). Different Gaussian peaks correspond to the charge responses for different numbers of PEs. By fitting the charge spectrum with a multi-Gaussian function, the gain is determined from the separation between adjacent peaks.
@@
Figure 9: Bias Between Gain Calibration Result and Mock Truth.
@@ 5.2 Internal Optical Crosstalk
When an electron undergoes an avalanche process in the avalanche region of the APD, the avalanche electron accelerates in the electric field and can emit photons. These photons have a probability of escaping from the APD and being detected by other APDs. This phenomenon is known as the optical crosstalk (OCT) effect in SiPMs [Ref44, Ref30]. Specifically, if an APD generates an OCT photon and this photon is detected by another APD within the same electronic channel, it is defined as internal optical crosstalk (IOCT). In contrast, if the APD detecting the OCT photon is in a different channel, it is referred to as external optical crosstalk (EOCT). Therefore, the OCT effect depends on the electric field intensity inside the APD, which is a function of the SiPM operating voltage. In TAO CD the operating voltage of the SiPM is stable.
There are two methods for calibrating the IOCT in the TAO CD: multiple photoelectrons (multi-PEs) hit analysis and generalized Poisson fitting. For multi-PEs hit analysis, the calibrated quantity $P_{\rm IOCT}$ represents the probability that a single electrical pulse generates IOCT pulses. In generalized Poisson fitting, the calibrated quantity $\lambda$ is the branching Poisson distribution parameter of the IOCT. It means that one pulse generates an average of $\lambda$ IOCT pulses following a Poisson distribution, which constitutes the first Poisson branch. Subsequently, each IOCT pulse in the first branch also generates average of $\lambda$ IOCT pulses, forming the second branch and subsequent branches follow similarly. The total distribution of these Poisson branches is branching Poisson distribution [Ref31]. The conversion relationship between these two quantities is given by Equation 5.2.
$P_{\rm IOCT}=1-exp(-\lambda)$ (5.2)
5.2.1 Multiple Photoelectron Hit Analysis
Hits occurring before the trigger time are generated by DN pulses and EOCT pulses induced by DN, which are referred to as equivalent dark noise (EDN) in this paper. These EDN pulses can induce IOCT. When IOCT is induced by these pulses, a multi-PEs hit composed of both the EDN and IOCT signals is invariably generated. In contrast, the probability of forming a multi-PE hit from EDN alone is minimal. As shown in Figure 10, it can be observed that multi-PEs hits are primarily contributed by the IOCT effect.
Figure 10: Charge Spectrum Distribution with and without IOCT. The red line represents a case with IOCT ($\lambda$ = 0.15), and the blue line represents a case without IOCT. It can be observed that the multi-PEs component is mainly contributed by IOCT.
Therefore, the multi-PEs region of the charge spectrum acquired in a dark environment is dominated by IOCT. Hence, the IOCT rate can be quantified using Equation 5.3, where $P_{\rm IOCT}$ is the probability of IOCT occurrence, $N_{\rm multPE}$ is the number of multi-PEs hits and $N_{\rm total}$ is the total number of hits.
@@ 5.2.1 Multiple Photoelectron Hit Analysis
Figure 11: Bias Between IOCT Calibration Result and Mock Truth. The parameter calibrated for IOCT corresponds to the probability of IOCT occurrence.
The IOCT calibration result is shown in Figure 11, the bias is 5.70% and the standard deviation is 2.73%. By analyzing the simulated data, which close IOCT and AP effect, the multi-PEs count rate is 0.74%. Therefore, a portion of the bias is attributable to the multi-PEs hits induced by EDN. The bias contribution is quantified by Equation 5.4, where $P_{\rm IOCT}$ is the probability of IOCT occurrence, $P_{\rm EDNM}$ is the probability that multi-PE hits are generated by EDN, $P_{\rm AP}$ is the probability of AP occurrence. The bias contributed by multi-PEs EDN is 4.57%. When the IOCT effect is disabled and 1% AP effect is added in the simulated data, the multi-PEs hit rate from 1 PE EDN hits is 0.23%, corresponding to a bias is 1.42%. It is important to note that since the IOCT calibration relies exclusively on DN hits, the limited statistics of the DN sample result in a relatively large standard deviation in the calibration results.
$\begin{split}Bias_{\rm EDNM}&=\frac{(1-P_{\rm IOCT})\cdot P_{\rm EDNM}}{P_{\rm IOCT}}\\ Bias_{\rm APM}&=\frac{(1-P_{\rm IOCT})\cdot P_{\rm AP}}{P_{\rm IOCT}}\end{split}$ (5.4)
@@ 5.2.2 Generalized Poisson Fitting
To eliminate the influence of multi-PEs EDN on the IOCT calibration, the charge spectrum can be fitted using a generalized Poisson distribution. As demonstrated in Vinogradov’s paper [Ref31], when $\mu$ primary pulses cause IOCT, the number of the first branch IOCT pulses generated by each primary pulse follows a Poisson distribution with a parameter $\lambda$. Each pulse in the first branch of IOCT can further generate the second branch of IOCT also following the same Poisson distribution; this pattern continues for subsequent branches. The total distribution of the number of pulses from the $\mu$ primary pulses and IOCT follows a the generalized Poisson distribution. Since IOCT pulses have no delay relative to the primary pulses, the distribution of the number of pulses corresponds to the number of PEs. The function of the generalized Poisson distribution function is shown in Equation 5.5, where k is the number of PE, $\mu$ is the number of primary pulse, and $\lambda$ is the branching Poisson distribution parameter of the IOCT.
For the EDN signal, $\mu=DCR^{*}\cdot t_{W}\cdot S$, where $DCR^{*}$ is equivalent DCR, $t_{W}$ is the average width of a single PE waveform, and S is the sensitive area of the SiPM in one channel. By replacing N in Equation 5.1 with GP in Equation 5.5 and fitting this function to the charge spectrum, the fitted value of $\lambda$ corresponds to the calibration result of IOCT.
The IOCT calibration result is shown in Figure 13, the bias is 1.40% and the standard deviation is 3.02%. According to the error propagation formula (Equation 5.2 to Equation 5.6), the bias caused by AP is 1.65%. The result indicates that the bias induced by multi-PEs EDN has been eliminated.
$Bias_{\rm APM,\lambda}=\frac{Bias_{\rm APM}}{1-P_{\rm IOCT}}$ (5.6)
Figure 12: Generalized Poisson Fit to the Charge Spectrum.
Figure 13: Bias Between IOCT Calibration Result and Mock Truth. The parameter calibrated for IOCT corresponds to the branch Poisson distribution parameter.
@@ 6 External Optical Crosstalk Calibration
The external optical crosstalk (EOCT) effect occurs when photons generated by avalanche electrons in APD escape from that APD and are detected by another APD in a different channel [Ref44, Ref30]. This paper presents a methodology for calibrating the EOCT rate and the distribution of EOCT emission angle in the TAO CD by switching the SiPM high voltage and using an LED external trigger to acquire data.
As shown in Equation 6.1, when all SiPMs in the TAO CD are operational, the electrical pulses generated within a SiPM include photoelectrons produced by the LED, DN, IOCT, AP, and EOCT from other SiPMs. Therefore, the charge $Q_{O}$ consists of contributions from the LED photon signal ($Q_{\rm LED}$), DN ($Q_{\rm DN}$), IOCT ($Q_{\rm IOCT}$), AP ($Q_{\rm AP}$) and EOCT ($Q_{\rm EOCT}$). When other SiPMs are turned off, the charge $Q_{C}$ only includes LED, DN, IOCT, and AP components. Thus, when acquiring "turn on" and "turn off" data over the same time duration, the difference between $Q_{O}$ and $Q_{C}$ corresponds to the contribution of EOCT. Here, the EOCT ratio is defined as $Q_{O}$ - $Q_{C}$ divided by $Q_{C}$.
$\begin{split}Q_{O}&=Q_{\rm LED}+Q_{\rm DN}+Q_{\rm IOCT}+Q_{\rm AP}+Q_{\rm EOCT}\\ Q_{C}&=Q_{\rm LED}+Q_{\rm DN}+Q_{\rm IOCT}+Q_{\rm AP}\\ Rate_{\rm EOCT}&=\frac{Q_{O}-Q_{C}}{Q_{C}}\end{split}$ (6.1)
@@ 6.1 External Optical Crosstalk Rate Calibration
Figure 14: EOCT Rate Calibration Result. The left figure presents the EOCT rate calibration value for each channel, with the definition of EOCT rate provided in Equation 6.1. The variation in EOCT rate across different layers is attributed to optical reflection at the SiPM surface. The middle and right figure shows the bias between the calibration value and the mock truth.
The bias and standard deviation are small because the calibrated EOCT rate (defined in Equation 6.1) is a directly observable quantity rather than a physical parameter such as the EOCT probability for each SiPM. The bias is quantified as the relative difference between the rates from Equation 6.1 and Equation 6.2 (In Equation 6.2, $Rate^{\rm truth}_{\rm EOCT}$ is the true EOCT ratio, $Q_{\rm EOCT}$ is the charge from EOCT, $Q_{C}$ is the charge measured when other SiPMs are turned off). Due to the LED sources can only be deployed along the central axis, the EOCT emission angle distribution must be calibrated to infer the response off the central axis.
$Rate^{\rm truth}_{\rm EOCT}=\frac{Q_{\rm EOCT}}{Q_{C}}$ (6.2)
@@ 6.2 External Optical Crosstalk Emission Angle Distribution Calibration
The external optical crosstalk emission angle is defined as the angle between the direction of EOCT photons escaping from a SiPM and the normal to the SiPM surface. In this study, the distribution of the SiPM EOCT emission angle is calibrated by using a reference SiPM to detect and quantify EOCT from other SiPMs positioned at different 5-degree angular intervals relative to the reference SiPM. The procedure is as follows: First, both the reference SiPM and the SiPMs within each 5-degree angular interval are turned on, and the total charge $Q_{O}$ on the reference SiPM is recorded. Next, only the reference SiPM is turned on, and data are acquired over the same duration to measure the total charge $Q_{C}$ on the reference SiPM. Finally, the EOCT response for each angular interval is obtained by subtracting $Q_{C}$ from $Q_{O}$.
The EOCT measurement response is the integrated response E from all SiPMs within each 5-degree angular interval. According to the mean value theorem for integral, the EOCT response $\overline{\epsilon(\bar{\theta})}$ in each 5-degree angular interval is given by Equation 6.3, where $\Delta\vec{S}$ is the area vector of each SiPM (with magnitude equals to the sensitive area and direction normal to the surface), $\vec{n}$ is the unit direction vector pointing from the emitting SiPM to the reference SiPM, $E(\theta)$ is the integrated EOCT response in each 5-degree angular interval, and $\epsilon(\theta)$ is the EOCT response at each angle.
$\begin{split}\overline{\epsilon(\bar{\theta})}&=\frac{E(\theta)}{\sum\vec{n}(\theta)\cdot\Delta\vec{S}(\theta)}\\ &=\frac{\sum\epsilon(\theta)\cdot\vec{n}(\theta)\cdot\Delta\vec{S}(\theta)}{\sum\vec{n}(\theta)\cdot\Delta\vec{S}(\theta)}\end{split}$ (6.3)
@@
Since the LED light field distribution may be non-uniform, this study simulated such a non-uniform LED light field distribution, as illustrated in Figure 16. The corresponding EOCT emission angle distribution calibration result is shown in the left part of Figure 17. This result indicates that the non-uniform LED light field distribution affects the EOCT emission angle distribution calibration.
Figure 16: LED Non-Uniform Light Field Distribution Configuration in Simulation. The distribution follows a Gaussian distribution (Gaus(90,30)), where the parameters correspond to the polar angle values in the spherical coordinate system.
In the TAO calibration system, the $^{68}{\rm Ge}$ source produces uniform scintillation light in the LS, allowing us to correct the non-uniformity of LED light field. Data maps for both the $^{68}{\rm Ge}$ source and the LED source are acquired with all SiPMs turned on. During the EOCT emission angle distribution calibration, the charge measured by the reference SiPM in the "turn on" data is multiplied by a correction factor $\alpha$ defined in Equation 6.4. The summation is performed over all SiPMs within the corresponding angular range, where $Q_{{}^{68}{\rm Ge}}$ represents the charge in the $^{68}{\rm Ge}$ data map, and $Q_{\rm LED}$ represents the charge in the LED data map.
@@ 6.3 Correct Dark Count Rate Base on the External Optical Crosstalk Rate Calibration Result
The DCR calibrated in Section 4.1 includes contributions from EOCT; therefore, a correction is necessary. The correction method is provided by Equation 6.5, where $\beta$ is the correction factor, $\delta_{\rm EOCT}$ is the average charge response to EOCT, $\delta_{\rm IOCT}$ is the average charge response to IOCT, Q is the charge measured by the SiPM from each effect, $Rate_{\rm EOCT}$ is the calibrated EOCT ratio, and $\lambda$ is the branching Poisson constant for IOCT. The values of $\delta_{\rm IOCT}$ can be derived from the mean value of the generalized Poisson distribution [Ref31]. Because the charge contribution of AP in the TAO SiPM mass test result is very low [Ref32], its contribution is neglected. Since $Rate_{\rm EOCT}$ is calibrated at the tile level, both channels within the same tile use the same $Rate_{\rm EOCT}$ value.
$\begin{split}DCR&=DCR^{*}\cdot\beta\\ Rate_{\rm EOCT}&=\frac{Q_{\rm EOCT}}{Q_{\rm LED}+Q_{\rm DN}+Q_{\rm IOCT}+Q_{\rm AP}}\\ &\approx\frac{Q_{\rm EOCT}}{Q_{\rm LED}+Q_{\rm DN}+Q_{\rm IOCT}}\\ &=\frac{\delta_{\rm EOCT}\cdot(Q_{\rm LED}+Q_{\rm DN})}{Q_{\rm LED}+Q_{\rm DN}+\delta_{\rm IOCT}\cdot(Q_{\rm LED}+Q_{\rm DN})}\\ \delta_{\rm EOCT}&=(1+\delta_{\rm IOCT})\cdot Rate_{\rm EOCT}\\ \delta_{\rm IOCT}&=\frac{\lambda}{1-\lambda}\\ \beta&=\frac{Q_{\rm DN}}{Q_{\rm DN}+Q_{\rm DN,\rm EOCT}}\\ &=\frac{Q_{\rm DN}}{Q_{\rm DN}+\delta_{\rm EOCT}\cdot Q_{\rm DN}}\\ &=\frac{1}{1+\delta_{\rm EOCT}}\\ &=\frac{1}{1+\frac{Rate_{\rm EOCT}}{1-\lambda}}\end{split}$ (6.5)
The result is shown in Figure 18. The bias is -0.40% and the standard deviation is 1.32%.

### v2

@@ 5.1 Gain
The charge generated by a single PE waveform follows a Gaussian distribution, and the individual photons detected by the APDs in the SiPM are independent. Consequently, the total charge resulting from the simultaneous detection of multiple PEs also follows a Gaussian distribution, and the overall distribution is a multi-Gaussian distribution (Equation 5.1 where f is the multi-Gaussian function, $N_{i}$ is the number of hits corresponding to each PE, $\mu_{i}$ is the mean value, and $\sigma_{i}$ is the standard deviation). Different Gaussian peaks correspond to the charge responses for different numbers of PEs. As shown in Figure 11, by fitting the charge spectrum with a multi-Gaussian function, the gain is determined from the separation between adjacent peaks.
$\begin{split}f=\sum_{i=1}^{n}N_{i}\cdot\text{Gaus}(\mu_{i},\sigma_{i})\end{split}$ (5.1)
@@
Figure 12: Gain Calibration Results. The left panel: comparison between the calibration results and the simulation truth in each channel. The right panel: 1D histogram filled with the each bin value from the left 2D histogram.
@@ 5.2 Internal Optical Crosstalk
There are two methods for calibrating the IOCT in the TAO CD: multiple PEs hit analysis and generalized Poisson fitting. For multiple PEs hit analysis, the calibrated quantity $P_{\rm IOCT}$ represents the probability that a single electrical pulse generates IOCT pulses. In generalized Poisson fitting, the calibrated quantity $\lambda_{\rm IOCT}$ is the Poisson parameter of the IOCT. According to the mathematical properties of the Generalized Poisson distribution [41], the conversion relationship between these two quantities is given by Equation 5.2.
$P_{\rm IOCT}=1-e^{-\lambda_{\rm IOCT}}$ (5.2)
5.2.1 Multiple Photoelectrons Hit Analysis
Hits occurring before the trigger time are generated by dark noise pulses and EOCT pulses induced by dark noise, which are referred to as equivalent dark noise in this paper. These equivalent dark noise pulses can induce IOCT. When IOCT is induced by these pulses, a multiple PEs hit composed of both the equivalent dark noise and IOCT signals is invariably generated. In contrast, the probability of forming a multiple PEs hit from equivalent dark noise alone is minimal. As shown in Figure 13, it can be observed that multiple PEs hits are primarily contributed by the IOCT effect.
Figure 13: Charge Spectrum Distribution with and without IOCT. The red line represents a case with IOCT ($\lambda_{\rm IOCT}$ = 0.15), and the blue line represents a case without IOCT. It can be observed that the multiple PEs component is mainly contributed by IOCT.
Therefore, the multiple PEs region of the charge spectrum acquired in a dark environment is dominated by IOCT. Hence, the IOCT rate can be quantified using Equation 5.3, where $P_{\rm IOCT}$ is the probability of IOCT occurrence, $N_{\rm multPE}$ is the number of multiple PEs hits and $N_{\rm total}$ is the total number of hits.
@@ 5.2.1 Multiple Photoelectrons Hit Analysis
Figure 14: IOCT Calibration Results Based on Multiple Photoelectrons Hit Analysis. The left panel: comparison between the calibration results and the simulation truth in each channel. The right panel: 1D histogram filled with the each bin value from the left 2D histogram.
The IOCT calibration result is shown in Figure 14, the bias is 5.70% and the standard deviation is 2.73%. The bias in the calibration from multiple PEs counting contributions caused by effects other than IOCT. By enabling and disabling various SiPM and electronics effects in the TAO simulation, we found that the multiple PEs hits excluding those from IOCT are dominated by two components: the multiple PEs hits generated by the equivalent dark noise itself, and contributions from the AP. The calibration bias introduced by the multiple PEs contribution of the equivalent dark noise itself is 4.57%, and the bias from the AP is 1.42%. It is clear that the IOCT calibration bias is dominated by the contribution of multiple PEs equivalent dark noise, which necessitates a dedicated correction scheme to mitigate and eliminate this bias.
@@ 5.2.2 Generalized Poisson Fitting
To eliminate the influence of multiple PEs equivalent dark noise on the IOCT calibration, the charge spectrum can be fitted using a generalized Poisson distribution. The total distribution of the number of pulses from the $\mu$ Poisson primary pulses and IOCT follows a the generalized Poisson distribution. Since IOCT pulses have no delay relative to the primary pulses, the distribution of the number of pulses corresponds to the number of PEs.
For the equivalent dark noise signal, $\mu=\text{DCR}^{*}\cdot t_{W}\cdot S$, where $\text{DCR}^{*}$ is equivalent DCR, $t_{W}$ is the average width of a single PE waveform, and S is the sensitive area of the SiPM in one channel. By replacing N in Equation 5.1 with GP in Equation 3.3 and fitting this function to the charge spectrum (as shown in Figure 15), the fitted value of $\lambda_{\rm IOCT}$ corresponds to the calibration result of IOCT.
The IOCT calibration result is shown in Figure 16, the bias is 1.40% and the standard deviation is 3.02%. This result indicates that the bias induced by multiple PEs equivalent dark noise has been eliminated.
Figure 15: Generalized Poisson Distribution Fit to the Charge Spectrum.
Figure 16: IOCT Calibration Results Based on Generalized Poisson Fitting. The left panel: comparison between the calibration results and the simulation truth in each channel. The right panel: 1D histogram filled with the each bin value from the left 2D histogram.
@@ 6 External Optical Crosstalk Calibration
This paper presents a methodology for calibrating the EOCT rate and the distribution of EOCT emission angle in the TAO CD by switching the SiPM high voltage and using an LED external trigger to acquire data.
As shown in Equation 6.1, when all SiPMs in the TAO CD are operational, the electrical pulses generated within a SiPM include PEs produced by the LED, dark noise, IOCT, AP, and EOCT from other SiPMs. Therefore, the charge $Q_{O}$ consists of contributions from the LED photon signal ($Q_{\rm LED}$), dark noise ($Q_{\rm DN}$), IOCT ($Q_{\rm IOCT}$), AP ($Q_{\rm AP}$) and EOCT ($Q_{\rm EOCT}$). When other SiPMs are turned off, the charge $Q_{C}$ only includes LED, dark noise, IOCT, and AP components. Thus, when acquiring "turn on" and "turn off" data over the same time duration, the difference between $Q_{O}$ and $Q_{C}$ corresponds to the contribution of EOCT. Here, the EOCT ratio is defined as $Q_{\rm EOCT}$ divided by $Q_{C}$.
$\begin{split}Q_{O}&=Q_{\rm LED}+Q_{\rm DN}+Q_{\rm IOCT}+Q_{\rm AP}+Q_{\rm EOCT}\\ Q_{C}&=Q_{\rm LED}+Q_{\rm DN}+Q_{\rm IOCT}+Q_{\rm AP}\\ \text{Rate}_{\rm EOCT}&=\frac{Q_{\rm EOCT}}{Q_{C}}\\ &=\frac{Q_{O}-Q_{C}}{Q_{C}}\end{split}$ (6.1)
@@ 6.1 External Optical Crosstalk Rate Calibration
Figure 17: EOCT calibration results based on the SiPM on-off switching method. The upper panel: SiPM-level EOCT calibration results. The middle panel: comparison between the calibration results and the simulation truth in each SiPM. The lower panel: 1D histogram filled with each bin value from the middle 2D histogram.
The bias and standard deviation are small because the calibrated EOCT rate (defined in Equation 6.1) is a directly observable quantity rather than a physical parameter such as the EOCT probability for each SiPM. The bias is quantified as the relative difference between the rates from Equation 6.1 and Equation 6.2 (In Equation 6.2, $\text{Rate}^{\rm truth}_{\rm EOCT}$ is the true EOCT ratio, $Q_{\rm EOCT}$ is the charge from EOCT, $Q_{C}$ is the charge measured when other SiPMs are turned off). Due to the LED sources can only be deployed along the central axis, the EOCT emission angle distribution must be calibrated to infer the response off the central axis.
$\text{Rate}^{\rm truth}_{\rm EOCT}=\frac{Q_{\rm EOCT}}{Q_{C}}$ (6.2)
@@ 6.2 External Optical Crosstalk Emission Angle Distribution Calibration
In this study, the distribution of the SiPM EOCT emission angle is calibrated by using a reference SiPM to detect and quantify EOCT from other SiPMs positioned at different 5-degree angular intervals relative to the reference SiPM. The procedure is as follows: First, both the reference SiPM and the SiPMs within each 5-degree angular interval are turned on, and the total charge $Q_{O}$ on the reference SiPM is recorded. Next, only the reference SiPM is turned on, and data are acquired over the same duration to measure the total charge $Q_{C}$ on the reference SiPM. Finally, the integrated EOCT response ($E(\theta)$) for each angular interval is obtained by subtracting $Q_{C}$ from $Q_{O}$.
According to the mean value theorem for integral, the EOCT response $\epsilon(\theta)$ in each 5-degree angular interval is given by Equation 6.3, where $\Delta\vec{S}$ is the area vector of each SiPM (with magnitude equals to the sensitive area and direction normal to the surface), $\vec{n}$ is the unit direction vector pointing from the emitting SiPM to the reference SiPM.
$\begin{split}\epsilon(\theta)&=\frac{E(\theta)}{\sum\vec{n}(\theta)\cdot\Delta\vec{S}(\theta)}\end{split}$ (6.3)
@@
Since the LED light field distribution may be non-uniform, this study simulated a non-uniform LED light field distribution, as illustrated in Figure 19. The corresponding EOCT emission angle distribution calibration result is shown in the left part of Figure 20. This result indicates that the non-uniform LED light field distribution affects the EOCT emission angle distribution calibration.
Figure 19: LED Non-Uniform Light Field Distribution Configuration in Simulation. The distribution follows a Gaussian distribution $\text{Gaus}(\theta,90^{\circ},30^{\circ})$, where the parameters correspond to the polar angle values in the spherical coordinate system.
In the TAO calibration system, the $^{68}{\rm Ge}$ source produces uniform scintillation light in the liquid scintillator, allowing us to correct the non-uniformity of LED light field. The charge response of each SiPM for both the $^{68}{\rm Ge}$ source and the LED source are acquired with all SiPMs turned on. During the EOCT emission angle distribution calibration, the charge measured by the reference SiPM in the "turn on" data is multiplied by a correction factor $\alpha$ defined in Equation 6.4. The summation is performed over all SiPMs within the corresponding angular range, $Q_{{\rm Ge}}$ denotes the SiPM charge response of to the $^{68}{\rm Ge}$ source, and $Q_{\rm LED}$ denotes the SiPM charge response of to the LED light source.
@@ 6.3 Correct Dark Count Rate Base on the External Optical Crosstalk Rate Calibration Result
The DCR calibrated in Section 4.1 includes contributions from EOCT, which is the equivalent dark count rate (EDCR). Therefore, a correction is necessary. The correction factor $\beta_{\rm DN}$:
$\begin{split}\text{DCR}&=\text{EDCR}\cdot\beta_{\rm DN}\\ \beta_{\rm DN}&=\frac{Q_{\rm DN}}{Q_{\rm DN}+Q_{\rm DN,\rm EOCT}}\end{split}$ (6.5)
This correction factor is characterized by the ratio of the dark noise charge response $Q_{\rm DN}$ to the total charge response from dark noise and dark noise-induced EOCT $Q_{\rm DN}+Q_{\rm DN,\rm EOCT}$.
Since the AP of the TAO SiPM is very low (<1%) [42], $\text{Rate}_{\rm EOCT}$ (from Equation 6.1) can be expressed as:
$\begin{split}\text{Rate}_{\rm EOCT}&=\frac{Q_{\rm EOCT}}{Q_{\rm LED}+Q_{\rm DN}+Q_{\rm IOCT}+Q_{\rm AP}}\\ &\approx\frac{Q_{\rm EOCT}}{Q_{\rm LED}+Q_{\rm DN}+Q_{\rm IOCT}}\\ &=\frac{\delta_{\rm EOCT}\cdot(Q_{\rm LED}+Q_{\rm DN})}{Q_{\rm LED}+Q_{\rm DN}+\delta_{\rm IOCT}\cdot(Q_{\rm LED}+Q_{\rm DN})}\end{split}$ (6.6)
where $\delta_{\text{EOCT}}$ denotes the mathematical expectation of EOCT generated by the avalanche signal from a photon hit or a dark noise event, and $\delta_{\text{IOCT}}$ denotes the mathematical expectation of IOCT generated by the avalanche signal from a photon hit or a dark noise event. For IOCT, the number of IOCT induced by an avalanche signal follows the Borel distribution. Based on the mathematical properties of the Borel distribution, $\delta_{\text{IOCT}}$ satisfies:
$\begin{split}\delta_{\rm IOCT}&=\frac{\lambda_{\rm IOCT}}{1-\lambda_{\rm IOCT}}\end{split}$ (6.7)
while $\delta_{\text{EOCT}}$ clearly satisfies:
$\begin{split}\delta_{\rm EOCT}&=\frac{Q_{\rm DN,\rm EOCT}}{Q_{\rm DN}}\end{split}$ (6.8)
According to Equation 6.5, Equation 6.6, Equation 6.7 and Equation 6.8, $\beta_{\rm DN}$ can be expressed as:
$\begin{split}\beta_{\rm DN}&=\frac{1}{1+\frac{\text{Rate}_{\rm EOCT}}{1-\lambda_{\rm IOCT}}}\end{split}$ (6.9)
The DCR result calibration after correction is shown in Figure 21. The bias is -0.40% and the standard deviation is 1.32%.
Figure 21: DCR calibration results corrected based on EOCT calibration results. The left panel: comparison between calibration results and the simulation truth in each channel. The right panel: 1D histogram filled with each bin value from the left 2D histogram.
7 Discussion
7.1 Impact of Temperature Fluctuation on the TAO SiPM Parameters
As reported in Ref. [9], the dependence of SiPM parameters on temperature and over voltage (OV) can be parameterized as:
$\begin{split}\lambda_{\rm IOCT}&=k_{\rm IOCT}\cdot\text{OV}\,,\qquad\lambda_{\rm EOCT}=k_{\rm EOCT}\cdot\text{OV}\,,\qquad\text{Gain}=k_{\text{Gain}}\cdot\text{OV}\,\\ \text{PDE}&=k_{\text{PDE}}\cdot(\text{OV}-\text{OV}_{\text{ref}})+\text{PDE}(\text{OV}_{\text{ref}})\\ \text{DCR}&=A_{\text{amp}}\cdot\text{OV}\cdot T^{3/2}\cdot e^{-E_{g}/kT}\ \end{split}$ (7.1)
According to the TAO SiPM burn-in test result [8], the relationship between the SiPM OV and temperature follows:
$\begin{split}\text{OV}=V_{\text{ex}}-V_{\text{bd}}=V_{\text{ex}}-(V_{\text{{bd}},{\text{ref}}}+k_{T}\cdot\Delta T)\end{split}$ (7.2)